# Articulation Work and Tinkering for Fairness in Machine Learning

### v1

@@ Abstract.
The field of fair AI aims to counter biased algorithms through computational modelling. However, it faces increasing criticism for perpetuating the use of overly technical and reductionist methods. As a result, novel approaches appear in the field to address more socially-oriented and interdisciplinary (SOI) perspectives on fair AI. In this paper, we take this dynamic as the starting point to study the tension between computer science (CS) and SOI research. By drawing on STS and CSCW theory, we position fair AI research as a matter of ‘organizational alignment’: what makes research ‘doable’ is the successful alignment of three levels of work organization (the social world, the laboratory and the experiment). Based on qualitative interviews with CS researchers, we analyze the tasks, resources, and actors required for doable research in the case of fair AI. We find that CS researchers engage with SOI to some extent, but organizational conditions, articulation work, and ambiguities of the social world constrain the doability of SOI research. Based on our findings, we identify and discuss problems for aligning CS and SOI as fair AI continues to evolve.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ACM Conference On Computer-Supported Cooperative Work And Social Computing; November 09-13, 2024; San José, Costa Ricaprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCIccs: Social and professional topics Socio-technical systemsccs: Computing methodologies Machine learning
@@ 1. Introduction
As of the present moment, fair AI finds itself at an intermediary phase of disciplinary evolvement (Heilbron, 2004; Merz and Sormani, 2016). Notably, fair AI is navigating internal dissent and external challenges about its future orientation (Weinberg, 2022; Miceli et al., 2022). This dynamic is marked by the interplay of two contrasting research paradigms: one rooted in computer science (CS), the origin discipline of fair AI, and another that is more socially-oriented and interdisciplinary (SOI) (Birhane et al., 2022; Benbouzid, 2023; Selbst et al., 2019). On one side, there is a defense of mathematically rigorous fairness approaches by CS researchers [^1] Dwork 2018. The Emerging Theory of Algorithmic Fairness. Microsoft Research (accessed via YouTube, 2024)., and an urgent call to not discredit them  [^2] Lidgett 2021. In Praise of Flawed Mathematical Models. ACM FAccT Conference (accessed via YouTube, 2024)., particularly in light of an increasing presence of AI technologies. On the other side, extensive critiques persistently challenge CS-oriented fair AI research. This criticism often stems from the perception that these approaches view solutions to discrimination and social inequality through a techno-optimistic lens, relying on computational methods without sufficient engagement with the reality of a socially stratified and diverse society (Davis et al., 2021). Scholars from various disciplines (Davis et al., 2021; Wachter et al., 2020; Alkhatib and Bernstein, 2019) have contributed to these critiques, contributing to the formation of the SOI paradigm. Boundaries are not always clear cut, including the case of the boundaries between CS and SOI paradigms and thus it should be noted that some prominent CS researchers have also contributed to the SOI paradigm by playing a key role in highlighting algorithmic discrimination (Buolamwini and Gebru, 2018; O’Neil, 2016) and offering self-critical perspectives in recent years (Narayanan, 2022; Ruggieri et al., 2023). Additionally, Abebe et al. (2020) highlights how CS can serve as a constructive ally in pursuing social change (an SOI oriented goal).
@@
In this paper, we suggest that the persisting gap between CS research and SOI critiques originates not (only) from epistemic and disciplinary tension, but is also rooted in organizational alignment. With this, we mean the alignment of research across three levels of work organization. According to theories in science and technology studies (STS) and computer-supported cooperative work (CSCW), research becomes only ‘doable’ when the three levels of work organization, e.g. the experiment, the laboratory and the social world, are successfully aligned (Fujimura, 1987; Strauss, 1985). As a result, researchers must spend time and effort in articulation work to achieve this alignment across all organizational levels. Further, researchers must gather the necessary resources, technologies, equipment and people, and align them, to do their research. Departing from this theoretical lens, and motivated by our own experiences of the tensions between CS and SOI, this study sets out to understand CS researchers’ perceptions and challenges of alignment with SOI in the case of fair AI.
@@
To address our inquiry, we carried out an interview study with a cohort of ten senior CS researchers based in different research centers and universities across Europe, who are actively working on fair AI research. Most of the interview participants were also collaborators of a European Union (EU)-funded project that focused on research and development of novel interdisciplinary methods for fair AI. We have noticed first-hand the prevalent role of CS in the produced outcomes of the project, thus additionally motivating our interest in how and to what degree CS researchers integrate the demands of SOI into their everyday research practices.
@@
The concept of articulation work also applies to other areas of CS research, as well as to research in general. However, studying the case of fair AI is particularly interesting given the increasing role of social world actors, such as regulatory and policy bodies. Our second contribution is thus to provide a deeper understanding of fair AI research as the emergent product of effectively aligning one’s research with these social world actors. We have personally experienced the effects that these actors can have on laboratory practices, and we consider it crucial to thoroughly examine the influence of such interactions.
Third, we are acutely aware that all too often, there is far too little time for a meaningful reflection on our daily work. The insight that our research is contingent on what is “doable” may initially seem quite straightforward. However, it also raises the question of how we can potentially expand the scope of doability. Connecting this to the identified gap between CS and SOI, we thirdly contribute with a reflection about how SOI can become a more doable problem.
@@
We first start with a presentation of the most important social world actors and their (partially ambiguous) demands. We find that there is an increased orientation towards the SOI paradigm for most actors, but also a lack of clarity about what doing SOI means. We then zoom in on the everyday practices on the level of the CS laboratory. We show that interview participants engage in SOI to some extent, but everyday work (e.g. writing proposals, securing funding and staff, reading and producing papers), technical requirements (e.g. obtaining data, and models) and temporal constraints challenge alignment with SOI, and restrict their ability to work beyond disciplinary expectations. We conclude our findings with the micro-strategies of the interview participants to integrate SOI into their everyday research practices.
@@
The rest of this paper is organized as follows: In Section 2, we give a brief overview of the evolvement of the two paradigms, CS and SOI, along with the development of fair AI research. We further outline important related work on the organizational challenges and constraints for fair AI development and research. In Section 3, we introduce the concept of articulation work, and explain how the doability of a research problem is contingent on the successful alignment across the three levels of work organization. In Section 4, we present our qualitative study and offer insights into the selection, academic background and scientific impact of our interview participants. We explain our approach to qualitative data coding and provide a detailed account of how we follow ethical standards in our interview study. In Section 5, we present a comprehensive analysis of our research outcomes. In Section 6, we discuss the role of facilitators and constraints in shaping the current state of fair AI research and its implications for making SOI-oriented fair AI research doable. We address the limitations of our research and propose future work in Section 7. Finally, Section 8 gives a brief reflection on potential directions for fair AI research.
@@ 2. Background and related work
Algorithmic-decision systems that draw on AI and machine learning (ML) techniques are proposed and deployed for a myriad of tasks across domains, some with high-stakes implications, e.g., healthcare treatment allocation, credit assessment, or job suitability. Their implementation is supported in great part by the promise “to bring greater discipline to decision-making” (Barocas et al., 2023). However, alongside the potential societal advantages attributed to the use of these systems, their deployment presents a challenge to society: they can perpetuate, and render invisible, structural forms of discrimination, such as sexism, racism, classism, and ageism. Extensively researched instances of algorithmic discrimination (Angwin et al., 2018; Buolamwini and Gebru, 2018; Lopez, 2021; Miceli et al., 2022) and media coverage of cases of algorithmic bias [^4] Dastin 2018. Amazon Scraps Secret AI Recruiting Tool that Showed Bias against Women. Reuters.  [^5] Guynn 2015. Google Photos labeled black people ’gorillas’. USA TODAY.  [^6] Angwin and Larson 2016. Machine Bias. ProPublica. have highlighted the imperative to detect and mitigate such harms. This has been paralleled by a growth in publications about AI ethical principles and guidelines (Jobin et al., 2019), the proposal of AI regulation across the globe (Commission, 2021; of Science and Technology Policy, 2022; Department of Industry, 2023; Yang, 2023), and the flourishing of related research communities across all disciplines. In this specific conjuncture, the field of fair AI has expanded across differing AI and computational sub-fields.
@@
Fair AI is currently situated at an “intermediate stage of disciplinary evolvement” (Heilbron, 2004; Merz and Sormani, 2016). In this stage, there is a growing importance of SOI, placing CS research on fair AI more and more at the intersection of these two research paradigms. As Benbouzid (2023) notes in regard to the increasing SOI demands for transparency, CS researchers “face two difficult and often distinct types of demands: first, for reliable computational techniques, and second, for transparency, given the constructed, politically situated nature of quantification operations”.
@@
Transparency is a characteristic that is asked of fair AI research, however within our context it is only one of the many indicators of the sociopolitical dimensions of this type of research. To expand on our introductory definition, we understand and use SOI as an umbrella term encompassing a broad range of research topics. SOI involves social scientist researchers as well as various societal actors, such as regulatory bodies, policymakers, ethicists, advocacy groups, and the broader community affected by AI technologies (Barocas et al., 2023). The gap between SOI and CS for fair AI also has led to the creation of and changes in already existing publication venues and conferences. The first Association for Computing Machinery [^7] Association for Computing Machinery (ACM) landing page(ACM)-sponsored conference organized by the FAT* network, now known as ACM Conference on Fairness, Accountability and Transparency (ACM FAccT), took place in 2018, leading to it become a renowned academic venue for innovative research on fairness, accountability, and transparency in ML with an interdisciplinary orientation [^8] ACM FAccT Conference: FAccT’22 Town Hall Meeting (accessed via YouTube, 2024).. The conference has strived to feature impactful work, and has given visibility to the ongoing criticism geared towards AI, such as the pursuit of technical progress at great material and environmental costs (Bender et al., 2021); inherent functional and external validity issues (Raji et al., 2022), and going further than that, its research values and foundations, as the work by Birhane et al. (Birhane et al., 2022) has shown.
@@
Alongside ACM FAccT, recent years have seen the emergence and establishment of other SOI-leaning venues, such as AAAI/ACM Conference on AI, Ethics, and Society (AIES), and ACM conference on Equity and Access in Algorithms, Mechanisms, and Optimization (EAAMO); venues which increasingly encourage interdisciplinary work, and are helmed by seemingly interdisciplinary leadership. Nevertheless, the gap between paradigms remains, and it can be seen in how the CS research community relates to and pursues different publishing venues. Given the popularity and interest of the fair AI topic, the last years have allowed for the emergence of new venues in the CS paradigm that explicitly seek to form a community supportive of primarily computationally rigorous research applied to “problems of pressing and anticipated societal concern” (e.g., Symposium on the Foundations of Responsible Computing). Similarly, there is now also the openness of traditional CS venues such as NeurIPS, the International Conference on Machine Learning (ICML), or the European Conference on Machine Learning and Principles and Practice of Knowledge Discovery in Databases (ECML-PKDD) to computational research on fair AI, via the main tracks, specially created ones or thematic workshops. A more recent emerging trend we have observed from leading and traditional CS venues (e.g., European Conference on Information Retrieval, and the International Joint Conference on Artificial Intelligence) has been the creation of special research tracks for AI that addresses societal concerns dedicated to work carried out in active collaboration with civil society stakeholders. Thus, when we refer to CS researchers aligning with SOI, we mean that they either engage in epistemic research topics that concern the SOI paradigms, engage in collaborations with SOI actors, or publish in SOI venues.
@@
Studies of the current tensions within fair AI have identified long-standing standards of practices, decision-making power imbalances, and fuzzy best practice guidelines to be all persistent challenges (Widder et al., 2022; Orr and Davis, 2020). Sloane and Zakrzewski (2022) also emphasize how ML systems, AI ethics, and social practices cannot be disentangled from their specific cultural and historical contexts. Multiple studies, (Holstein et al., 2019; Veale et al., 2018; Department of Industry, 2023; Heger et al., 2022), explore the gaps between research and the mundane, organizational work of CS researchers. Particularly, Holstein et al. (2019) and Veale et al. (2018), address existing challenges faced by practitioners in the private and public sector to develop and incorporate fairness notions in machine learning systems, while Deng et al. (2022) and Heger et al. (2022) examine the particulars of integrating fairness toolkits and documentation frameworks within organizational workflows.
@@ 3. Articulation work and the doability of research
Figure 1. Fujimura’s three levels of work organization (Fujimura, 1987), adapted for the case of fair AI. On the left side of the illustration, the three levels of work organization as per Fujimura’s (Fujimura, 1987) original conceptualization are depicted. To enable alignment, articulation work is needed, between levels. Further, the labelled lines (A & B) show how without alignment across all levels, doablity is not possible. On the right, we identify examples of actors and tasks that can be found across the three levels in line with our example.
@@
Fujimura (1987) examines three conditions of scientific research that facilitate or decrease articulation work. The leeway provided by available resources, and a clear division of labor facilitate articulating alignment. Conversely, uncertainty limits researchers’ ability to plan. Consequently, more articulation of work is then needed to conduct research spontaneously. As for the PhD student, constraints to performing their xAI user study could be the unavailability of courses during the semester, or the unwillingness of potential interview participants to remain throughout the length of the study. The student also needs to complete a convincing research proposal to get the support of their supervisor, regularly correspond with the hosting PI, and engage and survey relevant literature to support their findings. Suppose the hosting PI does not see value in performing the study, or the supervisor faces time constraints, articulation work for the student increases significantly. If alignment with important social world actors is not possible, the research will not be doable.
@@
To summarize, the doability of a research problem is contingent on the successful alignment across three levels of work organization, and shaped by facilitators and constraints, scientists’ actions and positions (Larregue, [n. d.]; Haraway, 1988; Harding, 2016). If an unexpected problem arrives, scientists engage in tinkering to maintain alignment with social world actors. Nevertheless, the impacts of organizational conditions on research are often overlooked, neglected or obscured (Lampland and Star, 2009; Nissenbaum, 2001).

### v2

@@ Abstract.
The field of fair AI aims to counter biased algorithms through computational modelling. However, it faces increasing criticism for perpetuating the use of overly technical and reductionist methods. As a result, novel approaches appear in the field to address more socially-oriented and interdisciplinary (SOI) perspectives on fair AI. In this paper, we take this dynamic as the starting point to study the tension between computer science (CS) and SOI research. By drawing on STS and CSCW theory, we position fair AI research as a matter of ‘organizational alignment’: what makes research ‘doable’ is the successful alignment of three levels of work organization (the social world, the laboratory, and the experiment). Based on qualitative interviews with CS researchers, we analyze the tasks, resources, and actors required for doable research in the case of fair AI. We find that CS researchers engage with SOI research to some extent, but organizational conditions, articulation work, and ambiguities of the social world constrain the doability of SOI research for them. Based on our findings, we identify and discuss problems for aligning CS and SOI as fair AI continues to evolve.
@@
copyright: rightsretainedjournal: PACMHCIjournalyear: 2024journalvolume: 8journalnumber: CSCW2article: 434publicationmonth: 11doi: 10.1145/3686973ccs: Human-centered computing Empirical studies in HCIccs: Social and professional topics Socio-technical systemsccs: Computing methodologies Machine learning
@@ 1. Introduction
As of the present moment, fair AI finds itself at an intermediary phase of disciplinary evolvement (Heilbron, 2004; Merz and Sormani, 2016). Notably, fair AI is navigating internal dissent and external challenges about its future orientation (Weinberg, 2022; Miceli et al., 2022). This dynamic is marked by the interplay of two contrasting research paradigms: one rooted in computer science (CS), the origin discipline of fair AI, and another one that is more socially-oriented and interdisciplinary (SOI) (e.g. (Birhane et al., 2022; Benbouzid, 2023; Selbst et al., 2019)). On one side, there is a defense of mathematically rigorous fairness approaches by CS researchers [^1] Dwork 2018. The Emerging Theory of Algorithmic Fairness. Microsoft Research (accessed via YouTube, 2024)., and an urgent call to not discredit them [^2] Lidgett 2021. In Praise of Flawed Mathematical Models. ACM FAccT Conference (accessed via YouTube, 2024)., particularly in light of the increasing presence of AI technologies. On the other side, extensive critiques persistently challenge CS-oriented fair AI research. This criticism often stems from the perception that these approaches view solutions to discrimination and social inequality through a techno-optimistic lens, relying on computational methods without sufficient engagement with the reality of a socially stratified and diverse society (Davis et al., 2021). Scholars from various disciplines (Davis et al., 2021; Wachter et al., 2020; Alkhatib and Bernstein, 2019) have voiced these critiques, contributing to the formation of the SOI paradigm. Boundaries are not always clear-cut, including the case of the boundaries between CS and SOI paradigms, and thus it should be noted that some prominent CS researchers have played a key role in highlighting algorithmic discrimination (Buolamwini and Gebru, 2018; O’Neil, 2016) and offering self-critical perspectives (Narayanan, 2022; Ruggieri et al., 2024). Additionally, Abebe et al. (2020) highlights how CS can serve as a constructive ally in pursuing social change (an SOI-oriented goal).
@@
In this paper, we suggest that the persisting gap between CS research and SOI critiques originates not (only) from epistemic and disciplinary tension, but is also rooted in organizational alignment. With this, we mean the alignment of research across three levels of work organization. According to theories in science and technology studies (STS) and computer-supported cooperative work (CSCW), research becomes only ‘doable’ when the three levels of work organization, e.g. the experiment, the laboratory, and the social world, are successfully aligned (Fujimura, 1987; Strauss, 1985). As a result, researchers must spend time and effort in articulation work to achieve this alignment across all organizational levels. Further, researchers must gather the necessary resources, technologies, equipment, and people, and align them, to do their research. Departing from this theoretical lens, and motivated by our own experiences of the tensions between the CS and SOI paradigms, this study sets out to understand CS researchers’ perceptions and challenges of alignment with SOI in the case of fair AI.
@@
To address our inquiry, we carried out an interview study with ten senior CS researchers based in different research centers and universities across Europe, who are actively working on fair AI research. Most of the interview participants were also collaborators of a European Union (EU)-funded project that focused on research and development of novel interdisciplinary methods for fair AI. We have noticed first-hand the prevalent role of CS in the produced outcomes of the project, thus additionally motivating our interest in how and to what degree CS researchers integrate the demands of SOI into their everyday research practices.
@@
The concept of articulation work also applies to other areas of CS research, as well as to research in general. However, studying the case of fair AI is particularly interesting given the increasing role of social world actors, such as regulatory and policy bodies. Our second contribution is thus to provide a deeper understanding of fair AI research as the emergent product of effectively aligning one’s research with social world actors. We have personally experienced the effects that these actors can have on laboratory practices, and we consider it crucial to thoroughly examine the influence of such interactions.
Third, we are acutely aware that all too often, there is far too little time for a meaningful reflection of our daily work. The insight that our research is contingent on what is “doable” may initially seem quite straightforward. However, it also raises the question of how we can potentially expand the scope of doability. Connecting this to the identified gap between CS and SOI, we thirdly contribute with a reflection on how SOI can become a more doable problem.
@@
We first start with a presentation of the most important social world actors and their (partially ambiguous) demands. We find that there is an increased orientation towards the SOI paradigm for most actors, but also a lack of clarity about what doing SOI research means. We then zoom in on the everyday practices on the level of the CS laboratory. We show that interview participants engage in SOI to some extent, but everyday work (e.g. writing proposals, securing funding and staff, reading and producing papers), technical requirements (e.g. obtaining data and models), and temporal constraints challenge alignment with SOI, and restrict their ability to work beyond disciplinary expectations. We conclude our findings with the micro-strategies of the interview participants to integrate SOI into their everyday research practices.
@@
The rest of this paper is organized as follows: In Section 2, we give a brief overview of the evolvement of the two paradigms, CS and SOI, along with the development of fair AI research. We further outline important related work on the organizational challenges and constraints for fair AI development and research. In Section 3, we introduce the concept of articulation work, and explain how the doability of a research problem is contingent on the successful alignment across the three levels of work organization. In Section 4, we present our qualitative study and offer insights into the selection, academic background, and scientific impact of our interview participants. We explain our approach to qualitative data coding and provide a detailed account of how we follow ethical standards in our interview study. In Section 5, we present a comprehensive analysis of our research outcomes. In Section 6, we discuss the role of facilitators and constraints in shaping the current state of fair AI research and its implications for making SOI-oriented fair AI research doable. We address the limitations of our research and propose future work in Section 7. Finally, Section 8 gives a brief reflection on potential directions for fair AI research.
@@ 2. Background and related work
Algorithmic-decision systems that draw on AI and machine learning (ML) techniques are proposed and deployed for a myriad of tasks across domains, some with high-stakes implications, e.g., healthcare treatment allocation, credit assessment, or job suitability. Their implementation is supported in great part by the promise “to bring greater discipline to decision-making” (Barocas et al., 2023). However, alongside the potential societal advantages attributed to the use of these systems, their deployment presents a challenge to society: they can perpetuate, and render invisible, structural forms of discrimination, such as sexism, racism, classism, and ageism. Extensively researched instances of algorithmic discrimination (Angwin et al., 2016; Buolamwini and Gebru, 2018; Lopez, 2021; Miceli et al., 2022) and media coverage of cases of algorithmic bias [^4] Dastin 2018. Amazon Scraps Secret AI Recruiting Tool that Showed Bias against Women. Reuters.  [^5] Guynn 2015. Google Photos labeled black people ’gorillas’. USA TODAY.  [^6] Angwin and Larson 2016. Machine Bias. ProPublica. have highlighted the imperative to detect and mitigate such harms. This has been paralleled by a growth in the number of publications about AI ethical principles and guidelines (Jobin et al., 2019), the proposal of AI regulation across the globe (Commission, 2021; of Science and Technology Policy, 2022; Department of Industry, 2023; Yang, 2023), and the flourishing of related research communities across all disciplines. In this specific conjuncture, the field of fair AI has also expanded across differing AI and computational sub-fields.
@@
Fair AI is currently situated at an “intermediate stage of disciplinary evolvement” (Heilbron, 2004; Merz and Sormani, 2016). In this stage, there is a growing importance of SOI research, placing CS research on fair AI more and more at the intersection of these two research paradigms. As Benbouzid (2023) notes in regard to the increasing SOI demands for transparency, CS researchers “face two difficult and often distinct types of demands: first, for reliable computational techniques, and second, for transparency, given the constructed, politically situated nature of quantification operations”.
@@
Transparency is a characteristic that is asked of fair AI research, however within our context, it is only one of the many indicators of the sociopolitical dimensions of this type of research. To expand on our introductory definition, we understand and use SOI as an umbrella term encompassing a broad range of research topics. SOI involves social scientist researchers as well as various societal actors, such as regulatory bodies, policymakers, ethicists, advocacy groups, and the broader community affected by AI technologies (Barocas et al., 2023). The gap between SOI and CS for fair AI also has led to the creation of, and changes in, already existing publication venues and conferences. The first Association for Computing Machinery [^7] Association for Computing Machinery (ACM) landing page(ACM)-sponsored conference organized by the FAT* network, now known as ACM Conference on Fairness, Accountability and Transparency (ACM FAccT), took place in 2018, leading to it becoming a renowned academic venue for innovative research on fairness, accountability, and transparency in ML with an interdisciplinary orientation [^8] ACM FAccT Conference: FAccT’22 Town Hall Meeting (accessed via YouTube, 2024).. The conference has striven to feature impactful work, and has given visibility to the ongoing criticism geared towards AI, such as the pursuit of technical progress at great material and environmental costs (Bender et al., 2021); inherent functional and external validity issues (Raji et al., 2022), and going further than that, its research values and foundations, as the work by Birhane et al. (2022) has shown.
@@
Alongside ACM FAccT, recent years have seen the emergence and establishment of other SOI-inclusive venues, such as AAAI/ACM Conference on AI, Ethics, and Society (AIES), and ACM conference on Equity and Access in Algorithms, Mechanisms, and Optimization (EAAMO); these venues increasingly encourage interdisciplinary work, and are helmed by seemingly interdisciplinary leadership. Nevertheless, the gap between paradigms remains, and it can be seen in how the CS research community relates to and pursues different publishing venues. Given the popularity and interest of the fair AI topic, the last years have allowed for the emergence of new venues in the CS paradigm that explicitly seek to form a community supportive of primarily computationally rigorous research applied to “problems of pressing and anticipated societal concern” (e.g., Symposium on the Foundations of Responsible Computing). Similarly, there is now also the openness of traditional CS venues such as NeurIPS, the International Conference on Machine Learning (ICML), or the European Conference on Machine Learning and Principles and Practice of Knowledge Discovery in Databases (ECML-PKDD) to computational research on fair AI, via the main tracks, specially created ones, or thematic workshops. A more recent emerging trend we have also observed from leading and traditional CS venues (e.g., European Conference on Information Retrieval, and the International Joint Conference on Artificial Intelligence) is the creation of special research tracks for AI that address societal concerns dedicated to work carried out in active collaboration with civil society stakeholders. Thus, when we refer to CS researchers aligning with SOI, we mean that they either engage in epistemic research topics that concern the SOI paradigms, engage in collaborations with SOI actors, or publish in SOI venues.
@@
Studies of the current tensions within fair AI have identified long-standing standards of practices, decision-making power imbalances, and fuzzy best practice guidelines to be all persistent challenges (Widder et al., 2022; Orr and Davis, 2020). Sloane and Zakrzewski (2022) also emphasize how ML systems, AI ethics, and social practices cannot be disentangled from their specific cultural and historical contexts. Multiple studies, (Holstein et al., 2019; Veale et al., 2018; Department of Industry, 2023; Heger et al., 2022), explore the gaps between research and the mundane, organizational work of CS researchers. Particularly, Holstein et al. (2019) and Veale et al. (2018), address existing challenges faced by practitioners in the private and public sectors to develop and incorporate fairness notions in machine learning systems, while Deng et al. (2022) and Heger et al. (2022) examine the particulars of integrating fairness toolkits and documentation frameworks within organizational workflows.
@@ 3. Articulation work and the doability of research
[…]
[Visualization of the levels of work organization represented with three polygons of different colors, side by side with a practical example.]On the left side of the illustration, there are three polygons organized vertically. At the top is the largest one, yellow-colored. It represents the Social World. The polygon in the middle is shorter, and green-colored, representing Laboratory Level. The one at the bottom, the smallest, and blue-colored, represents the Experiment Level. Here we have also three arrows illustrated, which represent research alignment for doable research problems. The first arrow (a) that only goes through the Social World is not a doable problem. The second arrow (b), which goes through the Social World and the Laboratory, also denotes a problem that is not doable. The third arrow, (c), goes through all three levels and thus denotes a doable problem. Further, we also include two ovals, one connects the Social World with the Laboratory to denote articulation work in between levels, and another one between Laboratory and Experiment. A bidirectional gray arrow unites this with the right side of the image. Here we have three-colored rectangles with rounded corners that contain details of our examples. In the Social World rectangle, which is yellow, we include the words: computer science, xAI conference chairs, psychology, and social science. In the Laboratory rectangle, which is green, the words are: Recruit interview partners, design interview study, attend UX courses, obtain recording device, purchase software for transcriptions. For the Experiment rectangle, which is blue, the words are: User study for framework elaboration, xAI framework elaboration.
Figure 1. Fujimura’s three levels of work organization (Fujimura, 1987), adapted for the case of fair AI. On the left side of the illustration, the three levels of work organization as per Fujimura’s (Fujimura, 1987) original conceptualization are depicted. To enable alignment, articulation work is needed, between levels. Further, the labeled lines (A & B) show how without alignment across all levels, doablity is not possible. On the right, we identify examples of actors and tasks that can be found across the three levels in line with our example.
@@
Fujimura (1987) examines three conditions of scientific research that facilitate or decrease articulation work. The leeway provided by available resources, and a clear division of labor facilitate articulating alignment. Conversely, uncertainty limits researchers’ ability to plan. Consequently, more articulation of work is then needed to conduct research spontaneously. As for the PhD student from our example, constraints to performing their xAI user study could be the unavailability of courses during the semester, or the unwillingness of potential interview participants to remain throughout the length of the study. The student also needs to complete a convincing research proposal to get the support of their supervisor, regularly correspond with the hosting PI, and engage and survey relevant literature to support their findings. However, in the event the hosting PI does not see value in performing the study, or the supervisor faces time constraints, articulation work for the student will increase significantly. If alignment with important social world actors is not possible, the research will not be doable.
@@
To summarize, the doability of a research problem is contingent on the successful alignment across three levels of work organization, and shaped by facilitators and constraints, scientists’ actions and positions (Larregue, 2018; Haraway, 1988; Harding, 2016). If an unexpected problem arises, scientists engage in tinkering to maintain alignment with social world actors. Nevertheless, the impacts of organizational conditions on research are often overlooked, neglected, or obscured (Lampland and Star, 2009; Nissenbaum, 2001).